# Akan Cinematic Emotions (ACE): A Multimodal Multi-party Dataset for Emotion Recognition in Movie Dialogues

### v2

@@ Abstract
In this paper, we introduce the Akan Conversation Emotion (ACE) dataset, the first multimodal emotion dialogue dataset for an African language, addressing the significant lack of resources for low-resource languages in emotion recognition research. ACE, developed for the Akan language, contains 385 emotion-labeled dialogues and 6,162 utterances across audio, visual, and textual modalities, along with word-level prosodic prominence annotations. The presence of prosodic labels in this dataset also makes it the first prosodically annotated African language dataset. We demonstrate the quality and utility of ACE through experiments using state-of-the-art emotion recognition methods, establishing solid baselines for future research. We hope ACE inspires further work on inclusive, linguistically and culturally diverse NLP resources.
Akan Cinematic Emotions (ACE): A Multimodal Multi-party Dataset
@@
David Sasu1, Zehui Wu2, Ziwei Gong2, Run Chen2, Pengyuan Shi2, Lin Ai2, Julia Hirschberg2, Natalie Schluter1 1IT University of Copenhagen, 2Columbia University
@@ 1 Introduction
To address this gap, we introduce the Akan Conversation Emotion (ACE) dataset, a multimodal emotion dialogue dataset for Akan, a major West African language spoken by about 20 million people Peterson et al. (n.d.). Akan is the most widely spoken language in Ghana, with around 80% of the population using it as a first or second language, and approximately 44% identifying as native Akan speakers. It is also natively spoken in parts of Ivory Coast and Togo. The language primarily comprises three main dialects: Asante, Akuapem, and Fante.
ACE contains 385 emotion-labeled dialogues from 21 Akan movies, covering diverse scenes and topics. It includes 6,162 utterances from 308 speakers (155 male, 153 female), ensuring a gender-balanced dataset. As a tonal language, Akan’s prosodic features are crucial for emotion recognition, so ACE includes word-level prosodic prominence annotations to support research on prosody in ERC. Our baseline experiments validate the dataset’s quality and utility for low-resource and cross-cultural emotion recognition research. Our main contributions are:
@@
    We introduce ACE, the first multimodal emotion dialogue dataset for an African language, enabling cross-cultural emotion research [^1] Our data and code are available at this anonymous repo..
@@
    We validate ACE through experiments with state-of-the-art ERC methods, establishing a strong baseline and detailed analysis.
@@
    We provide word-level prosodic prominence annotations, making ACE the first prosodically annotated dataset for an African language, facilitating research on prosody’s role in ERC and tonal language processing.
@@ 2.1 Related Datasets
ERC and Speech Emotion Recognition (SER) datasets are essential for developing models to understand and classify emotions in speech. These datasets can be grouped into three main categories. Table 1 compares ACE with other discussed datasets.
@@
Existing datasets lack resources for low-resource languages, such as Akan, and prosodic annotations critical for tonal languages. ACE fills this gap by offering a multi-modal resource with prosodic labels and conversational data, enabling robust SER for Akan and advancing cross-cultural SER research.
@@
ACE (Ours) Yes $a,v,t$ Yes 21 Movies No 7 308 Akan 6,162
@@ 2.2 Related Methods
3 ACE Dataset
We construct the ACE dataset by collecting and annotating dialogues from Akan-language movies, with examples illustrated in Figure 1. The dataset includes transcriptions with speaker identifications, emotion labels, and word-level prosodic prominence annotations. Table 1 compares ACE with other discussed datasets.
@@ 3.4 Emotion Annotation
The emotion demonstrated for each utterance within a dialogue was annotated using one of seven possible emotion labels: Sadness, Fear, Anger, Surprise, Disgust, Happy and Neutral. Six out of these emotions (i.e Sadness, Fear, Anger, Surprise, Happy and Disgust) were proposed by Paul Ekman (1992) as the six universal human emotions. Following previous works Poria et al. (2018); Busso et al. (2008); Gong et al. (2024), a neutral emotion label was added to identify utterances within the dataset that did not carry any pronounced emotional undertone.
@@ 3.5 Emotion Annotation Finalization
Table 2: General statistics of the ACE Dataset
@@ 3.6 Prosodic Prominence Annotation
At the time of writing, prosodic prominence labels from one annotator are complete and included in this paper. The second round of annotation is ongoing and will be incorporated into future versions of the dataset to enhance reliability through inter-annotator agreement analysis.
@@ 3.7 Dataset Statistics
Table 2 presents basic statistics of the Akan Cinematic Emotions (ACE) Dataset. It contains 385 dialogues, 4477 turns and 6162 utterances, which contain an average of 19 words. With respect to prosodic prominence, 37314 words were annotated to be prosodically prominent whereas 79991 words were annotated to be non-prominent.
@@
Table 3 illustrates the distribution of emotions in the ACE Dataset. Neutral emotion had the highest frequency, appearing in 2,941 instances, while Fear had the lowest frequency, occurring only 134 times.
@@
Table 3: Distribution of emotions in the ACE dataset
@@
The number of speakers in the ACE dataset is 308, of which 155 are men and 153 are women, as shown in Table 4.
@@
Table 4: Distribution of speakers in the ACE Dataset
@@ 4 Experiments and Analysis
We conduct a series of experiments to establish baseline performance for emotion recognition on ACE using unimodal and multimodal approaches. We first evaluate text, audio, and vision separately with state-of-the-art models, then explore modality combinations through feature concatenation and transformer-based fusion. These results serve as a foundation for future research on multimodal emotion recognition in Akan.
@@ 4.3 Audio Experiments
Spectrogram features are computed with 128 Mel-frequency bins, normalized, and truncated or padded to a maximum length of 1024 frames. These features are then used to fine-tune a pretrained Audio Spectrogram Transformer (AST) Gong et al. (2021) with an additional linear classifier layer. The model achieves a weighted F1 score of 47.89 and a macro F1 score of 23.36. We select AST due to its pretraining on a diverse auditory data, encompassing both human speech and non-human sounds, such as music and environmental noises. This broad training enables AST to effectively capture complex acoustic patterns, making it particularly well-suited for our ACE dataset, which consists of movie scenes containing a mix of dialogue, background music, and ambient sounds.
@@ 5 Conclusion and Future Directions
We introduce the Akan Conversation Emotion (ACE) dataset, the first multimodal emotion dialogue dataset for an African language, addressing the resource gap in ERC research for low-resource languages. ACE comprises 385 emotion-labeled dialogues and word-level prosodic prominence annotations, making it a valuable resource for cross-cultural emotion recognition and tonal language prosody research. Our experiments with state-of-the-art ERC methods validate ACE’s quality and establish a strong baseline for future research.
Looking ahead, we aim to expand to additional African languages and develop culturally adaptive ERC systems. Multimodal emotion recognition can be improved by speech enhancement techniques and pretraining models on African languages. Integrating vision-language models for scene descriptions can also provide richer context. Advanced fusion techniques like graph neural networks (GNNs) and hypergraphs may further refine cross-modal integration. We hope ACE inspires further research toward culturally adaptive, linguistically diverse NLP resources.
@@
While the Akan Cinematic Emotions (ACE) dataset represents a significant advancement in multimodal emotion recognition research, particularly for African languages, there are several limitations to acknowledge.
@@
Despite these limitations, we believe that ACE provides an essential foundation for advancing speech emotion recognition in low-resource languages and encourages further exploration in this area.
@@
The potential for misuse of the ACE dataset must be carefully acknowledged. While the dataset is intended for research purposes, deploying models trained on ACE in real-world applications without proper domain adaptation and validation could result in inaccurate emotion predictions, particularly in scenarios that deviate from cinematic dialogues. As such, researchers and practitioners should exercise caution when extending the use of this dataset to other applications.

### v3

# Akan Cinematic Emotions (AkaCE): A Multimodal Multi-party Dataset for Emotion Recognition in Movie Dialogues

@@ Abstract
In this paper, we introduce the Akan Conversation Emotion (AkaCE) dataset, the first multimodal emotion dialogue dataset for an African language, addressing the significant lack of resources for low-resource languages in emotion recognition research. AkaCE, developed for the Akan language, contains 385 emotion-labeled dialogues and 6,162 utterances across audio, visual, and textual modalities, along with word-level prosodic prominence annotations. The presence of prosodic labels in this dataset also makes it the first prosodically annotated African language dataset. We demonstrate the quality and utility of AkaCE through experiments using state-of-the-art emotion recognition methods, establishing solid baselines for future research. We hope AkaCE inspires further work on inclusive, linguistically and culturally diverse NLP resources.
Akan Cinematic Emotions (AkaCE): A Multimodal Multi-party Dataset
@@
David Sasu1, Zehui Wu2, Ziwei Gong2, Run Chen2, Pengyuan Shi2, Lin Ai2, Julia Hirschberg2, Natalie Schluter1,thanks: Currently at Apple. 1IT University of Copenhagen, 2Columbia University
@@ 1 Introduction
To address this gap, we introduce the Akan Conversation Emotion (AkaCE) dataset, a multimodal emotion dialogue dataset for Akan, a major West African language spoken by about 20 million people Peterson et al. (n.d.). Akan is the most widely spoken language in Ghana, with around 80% of the population using it as a first or second language, and approximately 44% identifying as native Akan speakers. It is also natively spoken in parts of Ivory Coast and Togo. The language primarily comprises three main dialects: Asante, Akuapem, and Fante.
AkaCE contains 385 emotion-labeled dialogues from 21 Akan movies, covering diverse scenes and topics. It includes 6,162 utterances from 308 speakers (155 male, 153 female), ensuring a gender-balanced dataset. As a tonal language, Akan’s prosodic features are crucial for emotion recognition, so AkaCE includes word-level prosodic prominence annotations to support research on prosody in ERC. Our baseline experiments validate the dataset’s quality and utility for low-resource and cross-cultural emotion recognition research. Our main contributions are:
@@
    We introduce AkaCE, the first multimodal emotion dialogue dataset for an African language, enabling cross-cultural emotion research [^1] Our data and code are available at this GitHub repo..
@@
    We validate AkaCE through experiments with state-of-the-art ERC methods, establishing a strong baseline and detailed analysis.
@@
    We provide word-level prosodic prominence annotations, making AkaCE the first prosodically annotated dataset for an African language, facilitating research on prosody’s role in ERC and tonal language processing.
@@ 2.1 Related Datasets
ERC and Speech Emotion Recognition (SER) datasets are essential for developing models to understand and classify emotions in speech. These datasets can be grouped into three main categories. Table 1 compares AkaCE with other discussed datasets.
@@
Existing datasets lack resources for low-resource languages, such as Akan, and prosodic annotations critical for tonal languages. AkaCE fills this gap by offering a multi-modal resource with prosodic labels and conversational data, enabling robust SER for Akan and advancing cross-cultural SER research.
@@
AkaCE (Ours) Yes $a,v,t$ Yes 21 Movies No 7 308 Akan 6,162
@@ 2.2 Related Methods
3 AkaCE Dataset
We construct the AkaCE dataset by collecting and annotating dialogues from Akan-language movies, with examples illustrated in Figure 1. The dataset includes transcriptions with speaker identifications, emotion labels, and word-level prosodic prominence annotations. Table 1 compares AkaCE with other discussed datasets.
@@ 3.4 Emotion Annotation
The emotion demonstrated for each utterance within a dialogue was annotated using one of seven possible emotion labels: Sadness, Fear, Anger, Surprise, Disgust, Happy and Neutral. Six out of these emotions (i.e Sadness, Fear, Anger, Surprise, Happy and Disgust) were proposed by Paul Ekman (1992) as the six universal human emotions. Following previous works Poria et al. (2018); Busso et al. (2008); Gong et al. (2024), a neutral emotion label was added to identify utterances within the dataset that did not carry any pronounced emotional undertone. This set of seven emotion labels is further supported by Ahmad et al. (2025), who employed the same categories in their study of cultural variation in emotion perception across 15 African languages.
@@ 3.5 Emotion Annotation Finalization
Table 2: General statistics of the AkaCE Dataset
@@ 3.6 Prosodic Prominence Annotation
An analysis of the inter-annotator agreement for prosodic annotation between the two annotators yielded an overall Fleiss’ Kappa statistic Fleiss (1971) of $k=\boldsymbol{1.0}$. The observed perfect agreement may be attributed to the tonal and expressive nature of Akan, which provides listeners with clear prosodic cues for identifying prominent words. In particular, prior work by Kügler and Genzel (2012) shows that Akan speakers signal pragmatic prominence through consistent acoustic patterns such as pitch lowering, even on high-toned syllables. This systematic use of prosody to express information structure likely contributes to the high salience of prominent words in speech, facilitating consistent annotation across raters.
@@ 3.7 Dataset Statistics
Table 2 presents basic statistics of the Akan Cinematic Emotions (AkaCE) Dataset. It contains 385 dialogues, 4477 turns and 6162 utterances, which contain an average of 19 words. With respect to prosodic prominence, 37314 words were annotated to be prosodically prominent whereas 79991 words were annotated to be non-prominent.
@@
Table 3 illustrates the distribution of emotions in the AkaCE Dataset. Neutral emotion had the highest frequency, appearing in 2,941 instances, while Fear had the lowest frequency, occurring only 134 times.
@@
Table 3: Distribution of emotions in the AkaCE dataset
@@
The number of speakers in the AkaCE dataset is 308, of which 155 are men and 153 are women, as shown in Table 4.
@@
Table 4: Distribution of speakers in the AkaCE Dataset
@@ 4 Experiments and Analysis
We conduct a series of experiments to establish baseline performance for emotion recognition on AkaCE using unimodal and multimodal approaches. We first evaluate text, audio, and vision separately with state-of-the-art models, then explore modality combinations through feature concatenation and transformer-based fusion. These results serve as a foundation for future research on multimodal emotion recognition in Akan.
@@ 4.3 Audio Experiments
Spectrogram features are computed with 128 Mel-frequency bins, normalized, and truncated or padded to a maximum length of 1024 frames. These features are then used to fine-tune a pretrained Audio Spectrogram Transformer (AST) Gong et al. (2021) with an additional linear classifier layer. The model achieves a weighted F1 score of 47.89 and a macro F1 score of 23.36. We select AST due to its pretraining on a diverse auditory data, encompassing both human speech and non-human sounds, such as music and environmental noises. This broad training enables AST to effectively capture complex acoustic patterns, making it particularly well-suited for our AkaCE dataset, which consists of movie scenes containing a mix of dialogue, background music, and ambient sounds.
@@ 4.5.2 Transformer Fusion
Emotion Class F1 Support
Neutral 68.57 409
Anger 51.50 176
Sadness 56.48 128
Joy 25.32 68
Surprise 15.21 26
Fear 0.08 16
Disgust 0.02 11
Table 10: Emotion-wide results of multimodal fusion experiments.
@@
Table 10 presents the per-emotion F1 scores and support. As expected, the model performs best on classes with higher frequency in the training set, such as Neutral, Anger, and Sadness, while performance on low-resource classes like Disgust and Fear remains limited. These findings underscore the importance of addressing class imbalance in multimodal emotion recognition, especially in low-resource language contexts such as Akan.
Model Complexity.
To contextualize the scale of our proposed transformer fusion system, we report in Table 11 the parameter counts for each of its core components. For the vision backbone, we use ResNet-18 (convolutional and fully connected layers), totalling approximately 11.7 million parameters. Text features are extracted using the Ghana-NLP abena-base-asante-twi-uncased model, a BERT-base variant with 12 transformer layers of 768 hidden units, contributing about 178 million parameters. Audio features are encoded with OpenAI’s Whisper-small model, which consists of 12-layer encoder-decoder stacks with 768-dimensional hidden representations, accounting for 244 million parameters. The transformer fusion module itself is composed of five cross-attention blocks, each using 768-dimensional queries, keys, and values with 12 attention heads, adding roughly 12 million parameters. Altogether, the complete model comprises approximately 450 million parameters, with the majority allocated to the unimodal backbones and only a small fraction to the fusion mechanism.
Component Configuration Parameters
ResNet-18 (Vision) Conv + FC layers 11.7M
BERT-base (Text) 12×768 (Abena-Twi) 178M
Whisper-small (Audio) 12×768 Encoder-Decoder 244M
Cross-Attention Blocks 5×(768-d, 12 heads) 12M
Total $\approx 450M$
Table 11: Model parameter breakdown across modalities.
@@ 5 Conclusion and Future Directions
We introduce the Akan Conversation Emotion (AkaCE) dataset, the first multimodal emotion dialogue dataset for an African language, addressing the resource gap in ERC research for low-resource languages. AkaCE comprises 385 emotion-labeled dialogues and word-level prosodic prominence annotations, making it a valuable resource for cross-cultural emotion recognition and tonal language prosody research. Our experiments with state-of-the-art ERC methods validate AkaCE’s quality and establish a strong baseline for future research.
Looking ahead, we aim to expand to additional African languages and develop culturally adaptive ERC systems. Multimodal emotion recognition can be improved by speech enhancement techniques and pretraining models on African languages. Integrating vision-language models for scene descriptions can also provide richer context. Advanced fusion techniques like graph neural networks (GNNs) and hypergraphs may further refine cross-modal integration. We hope AkaCE inspires further research toward culturally adaptive, linguistically diverse NLP resources.
Acknowledgements
This research is supported in part by the National Science Foundation via ARNI (The NSF AI Institute for Artificial and Natural Intelligence), under the Columbia 2025 Research Project ("Towards Safe, Robust, Interpretable Dialogue Agents for Democratized Medical Care"), and in part by the Defense Advanced Research Projects Agency (DARPA), via the CCU Program contract HR001122C0034. The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the National Science Foundation or the U.S. Government.
@@ Acknowledgements
While the Akan Cinematic Emotions (AkaCE) dataset represents a significant advancement in multimodal emotion recognition research, particularly for African languages, there are several limitations to acknowledge.
@@
Despite these limitations, we believe that AkaCE provides an essential foundation for advancing speech emotion recognition in low-resource languages and encourages further exploration in this area.
@@
The potential for misuse of the AkaCE dataset must be carefully acknowledged. While the dataset is intended for research purposes, deploying models trained on AkaCE in real-world applications without proper domain adaptation and validation could result in inaccurate emotion predictions, particularly in scenarios that deviate from cinematic dialogues. As such, researchers and practitioners should exercise caution when extending the use of this dataset to other applications.